# Empowering recommender systems using automatically generated Knowledge Graphs and Reinforcement Learning

### v1

@@ Abstract.
To this end, we present two knowledge graph-based approaches for personalized article recommendations for a set of customers of a large multinational financial services company. The first approach employs Reinforcement Learning (RL) and the second approach uses the XGBoost algorithm for recommending articles to the customers. Both approaches make use of a KG generated from both structured (tabular data) and unstructured data (a large body of text data).
@@
We also compared the above approaches with published algorithms for building recommender systems. Our proposed RL-based recommender system achieved 43.76% MAP (MAP@K=10). Our RL-based recommender system outperformed both the XGBoost-based approach and baseline model (Bayesian personalized ranking) by 13.38 and by 32.55 percentage points, respectively, delivering more accurate and personalized article recommendations. Importantly, our approach offers explainable results, promoting better decision-making. This study underscores the potential of combining advanced machine learning techniques with KG-driven insights to bolster experience in customer relationship management.
Knowledge Graph Embedding, TuckER, TransE, XAI
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: online and adaptive recommender systems; Aug 6–10, 2023; Long Beach, California, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Artificial Intelligence Machine Learningccs: Artificial Intelligence Recommender Systemsccs: Artificial Intelligence Reinforcement Learningccs: Artificial Intelligence XGBoostccs: Natural Language Processing Knowledge Graph
1. Introduction
@@
The increasing demand for personalized content has led to the development of recommendation systems that can effectively utilize structured information. Knowledge graphs (KGs) have emerged as a promising solution for this challenge, offering improved recommendation performance and explainability due to the inherent comprehensibility of relationships between entities (Nickel et al., 2011). A growing body of research is dedicated to exploring the potential of knowledge graph reasoning in personalized recommendation (Ai et al., 2018; Bordes et al., 2013; Grover and Leskovec, 2016; Wang et al., 2016). One line of research focuses on knowledge graph embedding models, such as TransE (Bordes et al., 2013) and node2vec (Grover and Leskovec, 2016), which align the knowledge graph in a regularized vector space, identifying the similarity between entities by calculating the distance between their representations (Zhang et al., 2016). However, purely KG embedding-based approaches struggle to uncover multi-hop relational paths, limiting the ability to capture complex relationships between entities. Another line of research investigates path-based recommendation techniques. Gao et al. (Gao et al., 2018) proposed the concept of meta-paths for reasoning over KGs. Although promising, this approach faces challenges when dealing with the numerous types of relations and entities present in large, real-world KGs, making it difficult to explore relationships between unconnected entities. Wang et al. (Wang et al., 2019b) developed a path embedding approach for recommendation over KGs that enumerates all qualified paths between every user-item pair, followed by training a sequential RNN model to predict ranking scores for the pairs. While this method improves recommendation performance, it is not feasible to explore all paths for every user-item pair in large-scale KGs due to computational limitations. Recent advances have focused on combining collaborative filtering (CF) with KG embedding techniques to enhance recommendation performance (Zhang et al., 2016; Ai et al., 2018). For example, Ai et al. (Ai et al., 2018) proposed a method that incorporated a soft matching algorithm to identify explanation paths between users and items. However, this strategy generates explanations post-hoc through empirical similarity matching between user and item embeddings, providing retrospective rationales for the chosen recommendations rather than deriving explanations from the reasoning process. We argue that an intelligent recommendation agent should explicitly reason over knowledge graphs for decision-making rather than simply embedding the graph as latent vectors for similarity matching. In this paper, we treat knowledge graphs as a flexible structure to maintain the agent’s knowledge about users, items, other entities, and their relationships. The agent initiates the process with a user and conducts explicit multi-step path reasoning over the graph, discovering suitable items for recommendation. This approach allows for the reasoning process to be easily interpreted, providing causal evidence for the recommended items. Our goal is not only to select a set of candidate items for recommendation but also to provide the corresponding reasoning paths as interpretable evidence for each recommendation. To address the limitations of previous work, we propose an approach that casts the recommendation problem as a deterministic Markov Decision Process (MDP) over the knowledge graph. We employ a Reinforcement Learning (RL) method, wherein an agent begins with a given user and learns to navigate to potential items of interest. The path in the KG then serves as an explanation for why the item should be recommended to the user. This approach presents three main challenges: measuring the correctness of an item for a user, efficiently exploring promising reasoning paths in the graph, and preserving the diversity of both items and paths during exploration. To tackle these challenges, we propose a KG-driven RL-based approach. The benefit of our approach is that it can also work when reviews or ratings of the items are not available and only click information is available to learn the user preferences.
@@
Our main contributions are as follows: (1) Automatic KG creation using structured and unstructured data. (2) Use of KG for building an XGBoost-based recommender system that can exploit click information. (3) Use of KG for building an RL-based recommender system that can exploit click information. (4) Explainability module that can explain the results.
@@
2. RELATED WORK
2.1. Collaborative Filtering
Collaborative Filtering (CF) has been a cornerstone in the development of recommender systems. Early approaches to CF focused on the user-item rating matrix and predicted ratings using user-based (Konstan et al., 1997; Resnick et al., 1994; Herlocker et al., 1999) or item-based (Linden et al., 2003; Sarwar et al., 2001) collaborative filtering methods. These approaches calculated similarities between users or items to generate recommendations.
As dimension reduction methods advanced, latent factor models, such as matrix factorization, gained widespread adoption in recommender systems. Prominent techniques include singular value decomposition (Koren et al., 2009), non-negative matrix factorization (Lee and Seung, 2000), and probabilistic matrix factorization (Mnih and Salakhutdinov, 2007). These methods essentially learn a latent factor representation for each user and item to calculate the matching score of user-item pairs.
In recent years, deep learning and neural models have further extended collaborative filtering, leading to two main sub-categories: similarity learning and representation learning. The similarity learning approach adopts relatively simple user/item embeddings (e.g., one-hot vectors) and learns a complex prediction network as a similarity function to compute user-item matching scores (He et al., 2017). In contrast, the representation learning approach focuses on learning richer user/item representations, while using a simple similarity function (e.g., inner product) for score matching (Zhang et al., 2017; Wang et al., 2019a).
@@ 2.1. Collaborative Filtering
In response to the challenges posed by the lack of interpretability in traditional collaborative filtering approaches, researchers have started to explore hybrid recommender systems that combine the benefits of CF methods with other techniques, such as knowledge graph-based methods (Guo et al., 2020; Zhang et al., 2016). These hybrid systems aim to improve the quality of recommendations while also providing more interpretable and explainable results.
@@
2.2. Knowledge Graph-driven Recommender Systems
Knowledge Graph-driven Recommender Systems (KGRS) have recently gained attention due to their ability to provide explainable and high-quality recommendations. Researchers have explored different ways to incorporate knowledge graph embeddings into recommender systems to improve recommendation performance and interpretability. One research direction focuses on leveraging knowledge graph embeddings as rich content information to enhance recommendation performance. For example, Zhang et al. (Zhang et al., 2016) utilized knowledge base embeddings to generate user and item representations for recommendation purposes. Huang et al. (Huang et al., 2018) employed memory networks over knowledge graph entity embeddings for recommendation. Wang et al. (Wang et al., 2018b) proposed a ripple network approach for embedding-guided multi-hop KG-based recommendation, which allows for the exploration of connections between entities in the knowledge graph. Another research direction aims to leverage the entity and path information in the knowledge graph to make explainable decisions. Ai et al. (Ai et al., 2018) incorporated the learning of knowledge graph embeddings for explainable recommendation, but their explanation paths are essentially post-hoc explanations, as they are generated by soft matching after the corresponding items have been chosen. Wang et al. (Wang et al., 2019b) proposed an RNN-based model to reason over KGs for recommendation. However, this approach requires enumerating all possible paths between each user-item pair for model training and prediction, which can be impractical for large-scale knowledge graphs.
@@ 2.2. Knowledge Graph-driven Recommender Systems
Some researchers have focused on leveraging the structural properties of knowledge graphs to improve recommendation performance. For instance, Wang et al. (Wang et al., 2019a) developed a graph attention network that incorporates both the relational information and entity features in a knowledge graph for recommendation. This approach allows for more accurate and context-aware recommendations by attending to the most relevant relations and entities for a given user-item pair.
In addition to using knowledge graph embeddings, researchers have also explored incorporating external knowledge sources and incorporating user-item interactions into the knowledge graph. Cao et al. (Cao et al., 2019) proposed a unified framework for incorporating user-item interactions and external knowledge sources into the knowledge graph, which improved the quality of recommendations by capturing the complex interplay between these elements.
Schlichtkrull et al. (Schlichtkrull et al., 2018) introduced a relational graph convolutional network (R-GCN) that learns embeddings for both entities and relations in a knowledge graph. This method can be used in a wide range of applications, including recommender systems, by exploiting the rich information present in the knowledge graph structure.
@@
2.3. Reinforcement Learning based Recommender Systems
@@
In non-KG settings, RL has been applied to various types of recommender systems, such as ads recommendation (Theocharous et al., 2015), news recommendation (Zheng et al., 2018), and post-hoc explainable recommendation (Wang et al., 2018a). These applications have demonstrated the potential of RL to adapt to changing user preferences and generate personalized recommendations based on user interactions.
In the context of knowledge graphs, researchers have primarily focused on utilizing RL for tasks such as question-answering (QA). For instance, Xiong et al. (Xiong et al., 2017) leveraged reinforcement learning for path-finding in knowledge graphs, while Das et al. (Das et al., 2017) proposed MINERVA which makes use of a KG and trains a model for question answering. Lin et al. (Lin et al., 2018) introduced RL-based models for KG question answering with reward shaping. These approaches formulate multi-hop reasoning as a sequential decision-making problem, taking advantage of the structure and information present in knowledge graphs.
@@ 3. METHODOLOGY
3.1. KG-Driven Reinforcement Learning based Recommender System
@@
The action space $A_{t}$ at state $s_{t}$ is defined as all possible emerging edges from an entity $e{t}$. Some nodes in the KG can have very large out-degree which can make it inefficient to maintain the large action space. Therefore, we perform an action-pruning step based on a scoring function $f((r,e)|u)$, which maps any relation to a real-valued score conditioned on a given user (Xian et al., 2019). There is a user-defined integer $\alpha$ that upper bounds the size of the action space. For our experiments, we set the value of $\alpha$ = 3.
@@
(3) $J(\theta)=\mathbb{E}_{e_{0}}\in u[\mathbb{E}_{a_{1},a_{2},...,a_{T}\sim\pi_{\theta}(a_{t}|s_{t})}[R_{T}]]$
To maximize the expected cumulative rewards, we use gradient ascent. The gradients are derived by the REINFORCE (Sutton and Barto, 2018), i.e.,
(4) $\bigtriangledown_{\theta}J(\theta)\approx\bigtriangledown_{\theta}\sum_{t}R_{T}log\pi_{\theta}(a_{t}|s_{t}).$
@@
3.2. KG-Driven XGBoost based Recommender System
XGBoost (eXtreme Gradient Boosting) (Chen and Guestrin, 2016) is an ensemble learning algorithm that has become a popular and effective method for a wide range of machine learning tasks, including classification, regression, and ranking. XGBoost builds a set of decision trees iteratively, using a gradient boosting approach to minimize a user-specified loss function.
For a dataset $D=\left\{(\textbf{x}_{i},y_{i})\right\}|(\textbf{x}_{i}\in\mathbb{R}^{m},y_{i}\in\mathbb{R})$ that has $n$ observations and $m$ features, the XGBoost model uses $Z$ additive functions for prediction (Chen and Guestrin, 2016).
(5) $\hat{y}_{i}=\sum_{z=1}^{Z}f_{k}(\textbf{x}_{i}),$
@@ 3.2. KG-Driven XGBoost based Recommender System
(6) $F=\left\{f(\textbf{x})=w_{q(x)}\right\}(q:\mathbb{R}^{m}\to T,w\in\mathbb{R}^{T}),$
@@
For our initial experiments, we implemented three rankers within the XGBoost model to predict the ranking of the articles for the users. These are XGBoost ranker (Chen and Guestrin, 2016), CatBoost ranker (Prokhorenkova et al., 2018), and LightGBM ranker (Ke et al., 2017). CatBoost (Prokhorenkova et al., 2018) is a recent library known for its efficacy in handling categorical features, which employs YetiRank (Gulin et al., 2011) as the loss function. LightGBM (Ke et al., 2017) handles categorical features and optimizes the LambdaRank loss. We trained XGBoost ranker  (Chen and Guestrin, 2016) with Rank Pairwise loss, utilizing one-hot encoding. During our initial experiments, the XGBoost ranker outperformed the other two rankers. Therefore, we selected the XGBoost ranker for our KG-driven XGBoost-based recommender system approach.
We used XGBoost ranker in combination with KGs generated from article text and the other article features to build the XGBoost-based recommender system. The KGs generated are then used as input to the TuckER and TransE to generate 300-dimensional KG embeddings. These embeddings along with the customer demographical data and article features are used to train the KG-driven XGBoost-based recommender system.
4. Experimental Setup

### v2

@@ Abstract
To this end, we present two knowledge graph-based approaches for educational article recommendations for a set of subscribers of a large multinational financial services company. The first approach employs Reinforcement Learning (RL) and the second approach uses the XGBoost algorithm for recommending articles to the subscribers. Both approaches make use of a KG generated from both structured (tabular data) and unstructured data (a large body of text data).
@@
We also compared the above approaches with published SOTA algorithms for building recommender systems. Our proposed RL-based recommender system achieved 43.76% MAP (MAP@K=10). Our RL-based recommender system outperformed both the XGBoost-based approach and baseline model (Bayesian personalized ranking) by 13.38 and by 32.55 percentage points, respectively, delivering more accurate and personalized educational article recommendations. Importantly, our approach offers explainable results, promoting better decision-making. This study underscores the potential of combining advanced machine learning techniques with KG-driven insights to bolster experience in customer relationship management.
Empowering recommender systems using automatically generated Knowledge Graphs and Reinforcement Learning
Ghanshyam Verma1, Simanta Sarkar1, Devishree Pillai1, Huan Chen1, John P. McCrae1, János A. Perge2, Shovon Sengupta2, Paul Buitelaar 1, 1Insight Research Ireland Centre for Data Analytics, Data Science Institute, University of Galway, Ireland 2Fidelity Investments, USA, Correspondence: ghanshyam.verma@insight-centre.org
1 Introduction
@@
The increasing demand for personalized content has led to the development of recommendation systems that can effectively utilize structured information. Knowledge graphs (KGs) have emerged as a promising solution for this challenge, offering improved recommendation performance and explainability due to the inherent comprehensibility of relationships between entities Nickel et al. (2011). A growing body of research is dedicated to exploring the potential of knowledge graph reasoning in personalized recommendation Ai et al. (2018); Bordes et al. (2013); Grover and Leskovec (2016); Wang et al. (2016). One line of research focuses on knowledge graph embedding models, such as TransE Bordes et al. (2013) and node2vec Grover and Leskovec (2016), which align the knowledge graph in a regularized vector space, identifying the similarity between entities by calculating the distance between their representations Zhang et al. (2016). However, purely KG embedding-based approaches struggle to uncover multi-hop relational paths, limiting the ability to capture complex relationships between entities. Another line of research investigates path-based recommendation techniques. Gao et al. Gao et al. (2018) proposed the concept of meta-paths for reasoning over KGs. Although promising, this approach faces challenges when dealing with the numerous types of relations and entities present in large, real-world KGs, making it difficult to explore relationships between unconnected entities. Wang et al. Wang et al. (2019b) developed a path embedding approach for recommendation over KGs that enumerates all qualified paths between every user-item pair, followed by training a sequential RNN model to predict ranking scores for the pairs. While this method improves recommendation performance, it is not feasible to explore all paths for every user-item pair in large-scale KGs due to computational limitations. Recent advances have focused on combining collaborative filtering (CF) with KG embedding techniques to enhance recommendation performance Zhang et al. (2016); Ai et al. (2018). For example, Ai et al. Ai et al. (2018) proposed a method that incorporated a soft matching algorithm to identify explanation paths between users and items. However, this strategy generates explanations post-hoc through empirical similarity matching between user and item embeddings, providing retrospective rationales for the chosen recommendations rather than deriving explanations from the reasoning process Xian et al. (2019a). We argue that an intelligent recommendation agent should explicitly reason over knowledge graphs for decision-making rather than simply embedding the graph as latent vectors for similarity matching. In this paper, we treat knowledge graphs as a flexible structure to maintain the agent’s knowledge about users, items, other entities, and their relationships. The agent initiates the process with a user and conducts explicit multi-step path reasoning over the graph, discovering suitable items for recommendation. This approach allows for the reasoning process to be easily interpreted, providing causal evidence for the recommended items. Our goal is not only to select a set of candidate items for recommendation but also to provide the corresponding reasoning paths as interpretable evidence for each recommendation. To address the limitations of previous work, we propose an approach that casts the recommendation problem as a deterministic Markov Decision Process (MDP) over the knowledge graph. We employ a Reinforcement Learning (RL) method, wherein an agent begins with a given user and learns to navigate to potential items of interest. The path in the KG then serves as an explanation for why the item should be recommended to the user. This approach presents three main challenges: measuring the correctness of an item for a user, efficiently exploring promising reasoning paths in the graph, and preserving the diversity of both items and paths during exploration. To tackle these challenges, we propose a KG-driven RL-based approach. The benefit of our approach is that it can also work when reviews or ratings of the items are not available and only click or other forms of response information is available to learn the user preferences.
@@
Our main contributions are as follows: (1) Automatic KG creation using structured and unstructured data. (2) Use of KG for building an XGBoost-based recommender system that can exploit click or other forms of response information. (3) Use of KG for building an RL-based recommender system that can exploit click/response information. (4) Explainability module that can explain the rationale behind the recommendations.
@@
2 RELATED WORK
2.1 Collaborative Filtering
Collaborative Filtering (CF) has been a cornerstone in the development of recommender systems. Early approaches to CF focused on the user-item rating matrix and predicted ratings using user-based Konstan et al. (1997); Resnick et al. (1994); Herlocker et al. (1999) or item-based Linden et al. (2003); Sarwar et al. (2001) collaborative filtering methods. These approaches calculated similarities between users or items to generate recommendations.
As dimension reduction methods advanced, latent factor models, such as matrix factorization, gained widespread adoption in recommender systems. Prominent techniques include singular value decomposition Koren et al. (2009), non-negative matrix factorization Lee and Seung (2000), and probabilistic matrix factorization Mnih and Salakhutdinov (2007). These methods essentially learn a latent factor representation for each user and item to calculate the matching score of user-item pairs.
In recent years, deep learning and neural models have further extended collaborative filtering, leading to two main sub-categories: similarity learning and representation learning. The similarity learning approach adopts relatively simple user/item embeddings (e.g., one-hot vectors) and learns a complex prediction network as a similarity function to compute user-item matching scores He et al. (2017). In contrast, the representation learning approach focuses on learning richer user/item representations, while using a simple similarity function (e.g., inner product) for score matching Zhang et al. (2017); Wang et al. (2019a).
@@ 2.1 Collaborative Filtering
In response to the challenges posed by the lack of interpretability in traditional collaborative filtering approaches, researchers have started to explore hybrid recommender systems that combine the benefits of CF methods with other techniques, such as knowledge graph-based methods Guo et al. (2020); Zhang et al. (2016). These hybrid systems aim to improve the quality of recommendations while also providing more interpretable and explainable results.
@@
2.2 Knowledge Graph-driven Recommender Systems
Knowledge Graph-driven Recommender Systems (KGRS) have recently gained attention due to their ability to provide explainable and high-quality recommendations. Researchers have explored different ways to incorporate knowledge graph embeddings into recommender systems to improve recommendation performance and interpretability. One research direction focuses on leveraging knowledge graph embeddings as rich content information to enhance recommendation performance. For example, Zhang et al. Zhang et al. (2016) utilized knowledge base embeddings to generate user and item representations for recommendation purposes. Huang et al. Huang et al. (2018) employed memory networks over knowledge graph entity embeddings for recommendation. Wang et al. Wang et al. (2018a) proposed a ripple network approach for embedding-guided multi-hop KG-based recommendation, which allows for the exploration of connections between entities in the knowledge graph. Another research direction aims to leverage the entity and path information in the knowledge graph to make explainable decisions. Ai et al. Ai et al. (2018) incorporated the learning of knowledge graph embeddings for explainable recommendation, but their explanation paths are essentially post-hoc explanations, as they are generated by soft matching after the corresponding items have been chosen. Wang et al. Wang et al. (2019b) proposed an RNN-based model to reason over KGs for recommendation. However, this approach requires enumerating all possible paths between each user-item pair for model training and prediction, which can be impractical for large-scale knowledge graphs.
@@ 2.2 Knowledge Graph-driven Recommender Systems
Some researchers have focused on leveraging the structural properties of knowledge graphs to improve recommendation performance. For instance, Wang et al. Wang et al. (2019a) developed a graph attention network that incorporates both the relational information and entity features in a knowledge graph for recommendation. This approach allows for more accurate and context-aware recommendations by attending to the most relevant relations and entities for a given user-item pair.
In addition to using knowledge graph embeddings, researchers have also explored incorporating external knowledge sources and incorporating user-item interactions into the knowledge graph. Cao et al. Cao et al. (2019) proposed a unified framework for incorporating user-item interactions and external knowledge sources into the knowledge graph, which improved the quality of recommendations by capturing the complex interplay between these elements.
Schlichtkrull et al. Schlichtkrull et al. (2018) introduced a relational graph convolutional network (R-GCN) that learns embeddings for both entities and relations in a knowledge graph. This method can be used in a wide range of applications, including recommender systems, by exploiting the rich information present in the knowledge graph structure.
@@
2.3 Reinforcement Learning based Recommender Systems
@@
In non-KG settings, RL has been applied to various types of recommender systems, such as ads recommendation Theocharous et al. (2015), news recommendation Zheng et al. (2018), and post-hoc explainable recommendation Wang et al. (2018b). These applications have demonstrated the potential of RL to adapt to changing user preferences and generate personalized recommendations based on user interactions.
In the context of knowledge graphs, researchers have primarily focused on utilizing RL for tasks such as question-answering (QA). For instance, Xiong et al. Xiong et al. (2017) leveraged reinforcement learning for path-finding in knowledge graphs, while Das et al. Das et al. (2017) proposed MINERVA which makes use of a KG and trains a model for question answering. Lin et al. Lin et al. (2018) introduced RL-based models for KG question answering with reward shaping. These approaches formulate multi-hop reasoning as a sequential decision-making problem, taking advantage of the structure and information present in knowledge graphs.
@@ 3 METHODOLOGY
3.1 KG-Driven Reinforcement Learning based Recommender System
@@
The action space $A_{t}$ at state $s_{t}$ is defined as all possible emerging edges from an entity $e{t}$. Some nodes in the KG can have very large out-degree which can make it inefficient to maintain the large action space. Therefore, we perform an action-pruning step based on a scoring function $f((r,e)|u)$, which maps any relation to a real-valued score conditioned on a given user Xian et al. (2019b). There is a user-defined integer $\alpha$ that upper bounds the size of the action space. For our experiments, we set the value of $\alpha$ = 3.
@@
$J(\theta)=\mathbb{E}_{e_{0}}\in u[\mathbb{E}_{a_{1},a_{2},...,a_{T}\sim\pi_{\theta}(a_{t}|s_{t})}[R_{T}]]$ (3)
To maximize the expected cumulative rewards, we use gradient ascent. The gradients are derived by the REINFORCE Sutton and Barto (2018), i.e.,
$\bigtriangledown_{\theta}J(\theta)\approx\bigtriangledown_{\theta}\sum_{t}R_{T}log\pi_{\theta}(a_{t}|s_{t}).$ (4)
@@
3.2 KG-Driven XGBoost based Recommender System
XGBoost (eXtreme Gradient Boosting) Chen and Guestrin (2016) is an ensemble learning algorithm that has become a popular and effective method for a wide range of machine learning tasks, including classification, regression, and ranking. XGBoost builds a set of decision trees iteratively, using a gradient boosting approach to minimize a user-specified loss function.
For a dataset $D=\left\{(\textbf{x}_{i},y_{i})\right\}|(\textbf{x}_{i}\in\mathbb{R}^{m},y_{i}\in\mathbb{R})$ that has $n$ observations and $m$ features, the XGBoost model uses $Z$ additive functions for prediction Chen and Guestrin (2016).
$\hat{y}_{i}=\sum_{z=1}^{Z}f_{k}(\textbf{x}_{i}),$ (5)
@@ 3.2 KG-Driven XGBoost based Recommender System
$F=\left\{f(\textbf{x})=w_{q(x)}\right\}(q:\mathbb{R}^{m}\to T,w\in\mathbb{R}^{T}),$ (6)
@@
For our initial experiments, we implemented three rankers within the XGBoost model to predict the ranking of the articles for the users. These are XGBoost ranker Chen and Guestrin (2016), CatBoost ranker Prokhorenkova et al. (2018), and LightGBM ranker Ke et al. (2017). CatBoost Prokhorenkova et al. (2018) is a recent library known for its efficacy in handling categorical features, which employs YetiRank Gulin et al. (2011) as the loss function. LightGBM Ke et al. (2017) handles categorical features and optimizes the LambdaRank loss. We trained XGBoost ranker  Chen and Guestrin (2016) with Rank Pairwise loss, utilizing one-hot encoding. During our initial experiments, the XGBoost ranker outperformed the other two rankers. Therefore, we selected the XGBoost ranker for our KG-driven XGBoost-based recommender system approach.
We used XGBoost ranker in combination with KGs generated from article text and the other article features to build the XGBoost-based recommender system. The KGs generated are then used as input to the TuckER and TransE to generate 300-dimensional KG embeddings. These embeddings along with the subscriber demographical data and educational article features are used to train the KG-driven XGBoost-based recommender system.
4 Experimental Setup